Circular No. 7175

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 K2 (FERRIS) LONEOS reports the discovery of a comet by William D. Ferris on

CCD frames taken with the 0.59-m LONEOS Schmidt telescope:

1999 UT  α_{2000}  δ_{2000}  m_{1}
May 19.36625  20^{h}48^{m}11.^{s}34  −7^{o}56′29′′.5  18.1

Measurer B. Koehn notes that the comet showed a well-condensed nucleus, a coma of diameter ∼ 15′′, and a faint tail ∼ 20′′ long in p.a. 225^{o} on May 19. Additional astrometry appears on MPEC 1999-K22. May 22 observations by J. Tichá and M. Tichý (Kleť) show the comet as diffuse with a 12′′ coma; observations on the same night by L. Kornoš and P. Koleny (Modra) also show a coma. C. W. Hergenrother and A. E. Gleason (Catalina 1.5-m reflector) report a 20′′ coma and a 20′′ tail in p.a. 230^{o}. The following very preliminary parabolic orbital elements are by B. G. Marsden:

T = 1999 Oct. 30.589 TT  ω = 28.^{o}089
Ω = 300.677 }2000.0
q = 5.16257 AU  i = 84.306

COMET C/1999 K3 (LINEAR) M. Elowitz, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the discovery of an apparently cometary object in LINEAR data, as follows:

1999 UT  α_{2000}  δ_{2000}  m_{1}
May 20.26972  19^{h}02^{m}40.^{s}60  +9^{o}02′47′′.8  18.7

Following of this object on The NEO Confirmation Page, numerous observers confirmed cometary appearance, and the additional astrometry and orbit are given on MPEC 1999-K23. Around May 22.0 UT, L. Šarounová (Ondřejov) reports coma diameters ∼ 20′′ and 15′′; Tichá and Tichý report a 10′′ coma and a wide tail in p.a. 245^{o}; and coma was also noted by Kornoš and Koleny. Preliminary parabolic elements by Marsden:

T = 1999 Feb. 26.754 TT  ω = 340.^{o}458
Ω = 266.798 }2000.0
q = 1.91752 AU  i = 92.027
1999 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green